### v1

@@ 3 Moment Estimates
$=$ $\frac{1}{\epsilon^{2}}\int_{s}^{t}\int_{0}^{u}z^{\epsilon,r}_{u}dW_{r}du+\frac{1}{\epsilon^{2}}\int_{s}^{t}\int_{0}^{u}z^{\epsilon,r}_{u}F(X^{\epsilon}_{r})drdu$
$=$ $\Big[\frac{1}{\epsilon^{2}}\int_{0}^{s}\int_{s}^{t}z^{\epsilon,r}_{u}dudW_{r}+\frac{1}{\epsilon^{2}}\int_{s}^{t}\int_{r}^{t}z^{\epsilon,r}_{u}dudW_{r}\Big]$
$+\Big[\frac{1}{\epsilon^{2}}\int_{0}^{s}\int_{s}^{t}z^{\epsilon,r}_{u}F(X^{\epsilon}_{r})dudr+\frac{1}{\epsilon^{2}}\int_{s}^{t}\int_{r}^{t}z^{\epsilon,r}_{u}F(X^{\epsilon}_{r})dudr\Big].$
Now we deal with four terms above respectively. By Burkholder–Davis–Gundy inequality,
$\mathbb{E}\Big|\int_{0}^{s}\int_{s}^{t}z^{\epsilon,r}_{u}dudW_{r}\Big|^{p}\lesssim\mathbb{E}\Big(\int_{0}^{s}\Big|\int_{s}^{t}z^{\epsilon,r}_{u}du\Big|^{2}dr\Big)^{p/2}.$ (3.11)
From (3.4) and Hölder inequality
$\Big|\int_{s}^{t}z^{\epsilon,r}_{u}du\Big|^{2}$
$\leq$ $\Big|\int_{s}^{t}e^{-\epsilon^{-2}(u-r)}du\Big|^{2}$
$\leq$ $(t-s)\int_{s}^{t}e^{-2\epsilon^{-2}(u-r)}du.$
Therefore,
$\int_{0}^{s}\Big|\int_{s}^{t}z^{\epsilon,r}_{u}du\Big|^{2}dr\lesssim(t-s)\int_{0}^{s}\int_{s}^{t}e^{-2\epsilon^{-2}(u-r)}dudr\lesssim\epsilon^{4}(t-s).$ (3.12)
Substituting the above inequality into (3.11),
$\mathbb{E}\Big|\int_{0}^{s}\int_{s}^{t}z^{\epsilon,r}_{u}dudW_{r}\Big|^{p}\lesssim\epsilon^{2p}|t-s|^{p/2}.$ (3.13)
By Burkholder–Davis–Gundy inequality and (3.4),
$\mathbb{E}\Big|\int_{s}^{t}\int_{r}^{t}z^{\epsilon,r}_{u}dudW_{r}\Big|^{p}$ (3.14)
$\lesssim$ $\mathbb{E}\Big(\int_{s}^{t}\Big|\int_{r}^{t}z^{\epsilon,r}_{u}du\Big|^{2}\Big)^{p/2}$
$\leq$ $\mathbb{E}\Big(\int_{s}^{t}\Big|\int_{r}^{t}e^{-\epsilon^{-2}(u-r)}du\Big|^{2}dr\Big)^{p/2}$
$\leq$ $\Big(\int_{s}^{t}\epsilon^{4}dr\Big)^{p/2}$
$\leq$ $\epsilon^{2p}(t-s)^{p/2}.$
On the other hand, by ($\mathbf{A4}$), Hölder inequality and (3.12),
$\mathbb{E}\Big|\int_{0}^{s}\int_{s}^{t}z^{\epsilon,r}_{u}F(X^{\epsilon}_{r})dudr\Big|^{p}$ (3.15)
$\lesssim$ $\mathbb{E}\Big(\int_{0}^{s}\int_{s}^{t}|z^{\epsilon,r}_{u}|dudr\Big)^{p}$
$=$ $\mathbb{E}\Big\{\Big[\Big(\int_{0}^{s}\int_{s}^{t}|z^{\epsilon,r}_{u}|dudr\Big)^{2}\Big]^{p/2}\Big\}$
$\lesssim$ $\mathbb{E}\Big\{\Big[\int_{0}^{s}\Big|\int_{s}^{t}z^{\epsilon,r}_{u}du\Big|^{2}dr\Big]^{p/2}\Big\}$
$\lesssim$ $\epsilon^{2p}|t-s|^{p/2}.$
Lastly,
$\mathbb{E}\Big|\int_{s}^{t}\int_{r}^{t}z^{\epsilon,r}_{u}F(X^{\epsilon}_{r})dudr\Big|^{p}$ (3.16)
$\lesssim$ $\mathbb{E}\Big|\int_{s}^{t}\int_{r}^{t}e^{-\epsilon^{-2}(u-r)}dudr\Big|^{p}$
$\leq$ $\Big|\int_{s}^{t}\epsilon^{2}dr\Big|^{p}$
$\leq$ $\epsilon^{2p}(t-s)^{p}.$
Substituting (3.13)–(3.16) into (3.10),
$\mathbb{E}|X^{\epsilon}_{t}-X^{\epsilon}_{s}|^{p}\lesssim|t-s|^{p/2}.$ (3.17)
We turn to prove (ii). By triangular inequality, it is sufficient to consider the $(i,j)$-entry $\int_{s}^{t}X^{\epsilon,i}_{s,u}dX^{\epsilon,j}_{u}$ of $\int_{s}^{t}X^{\epsilon}_{s,u}\otimes dX^{\epsilon}_{u}.$ But from (3.3),
$dX^{\epsilon,j}_{t}=\frac{1}{\epsilon^{2}}\sum_{k}\int_{0}^{t}z^{\epsilon,r}_{t}(j,k)dB^{k}_{r}dt+\frac{1}{\epsilon^{2}}\sum_{k}\int_{0}^{t}z^{\epsilon,r}_{t}(j,k)F_{k}(X^{\epsilon}_{r})drdt,$
where $z^{\epsilon,r}_{t}(j,k)$ is the ($j,k$)-entry of $z^{\epsilon,r}_{t}$, so
$\int_{s}^{t}X^{\epsilon,i}_{s,u}dX^{\epsilon,j}_{u}=\frac{1}{\epsilon^{2}}\sum_{k}\int_{s}^{t}\int_{0}^{u}X^{\epsilon,i}_{s,u}z^{\epsilon,r}_{u}(j,k)dB^{k}_{r}du+\frac{1}{\epsilon^{2}}\sum_{k}\int_{s}^{t}\int_{0}^{u}X^{\epsilon,i}_{s,u}z^{\epsilon,r}_{u}(j,k)F_{k}(X^{\epsilon}_{r})drdu.$
Without loss of generality, we assume that the system is one-dimensional, that is, we write
$\int_{s}^{t}X^{\epsilon}_{s,u}dX^{\epsilon}_{u}=\frac{1}{\epsilon^{2}}\int_{s}^{t}\int_{0}^{u}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}dW_{r}du+\frac{1}{\epsilon^{2}}\int_{s}^{t}\int_{0}^{u}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}F(X^{\epsilon}_{r})drdu.$ (3.18)
By stochastic Fubini theorem,
$\int_{s}^{t}X^{\epsilon}_{s,u}dX^{\epsilon}_{u}$ (3.19)
$=$ $\Big[\frac{1}{\epsilon^{2}}\int_{0}^{s}\int_{s}^{t}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}dudW_{r}+\frac{1}{\epsilon^{2}}\int_{s}^{t}\int_{r}^{t}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}dudW_{r}\Big]$
$+\Big[\frac{1}{\epsilon^{2}}\int_{0}^{s}\int_{s}^{t}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}F(X^{\epsilon}_{r})dudr+\frac{1}{\epsilon^{2}}\int_{s}^{t}\int_{r}^{t}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}F(X^{\epsilon}_{r})dudr\Big]$
$=:$ $I+J.$
We firstly estimate $I$. Using Burkholder–Davis–Gundy inequality,
$\mathbb{E}|I|^{p}$ (3.20)
$\lesssim$ $\frac{1}{\epsilon^{2p}}\mathbb{E}\Big|\int_{0}^{s}\int_{s}^{t}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}dudW_{r}\Big|^{p}+\frac{1}{\epsilon^{2p}}\mathbb{E}\Big|\int_{s}^{t}\int_{r}^{t}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}dudW_{r}\Big|^{p}$
$\lesssim$ $\frac{1}{\epsilon^{2p}}\mathbb{E}\Big(\int_{0}^{s}\Big|\int_{s}^{t}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}du\Big|^{2}dr\Big)^{p/2}+\frac{1}{\epsilon^{2p}}\mathbb{E}\Big(\int_{s}^{t}\Big|\int_{r}^{t}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}du\Big|^{2}dr\Big)^{p/2}.$
Applying Lemma 3.2 to $G(r)=\int_{s}^{t}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}du$ and $q=p/2,$
$\Big\{\mathbb{E}\Big(\int_{0}^{s}\Big|\int_{s}^{t}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}du\Big|^{2}dr\Big)^{p/2}\Big\}^{2/p}\leq\int_{0}^{s}\Big(\mathbb{E}\Big|\int_{s}^{t}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}du\Big|^{p}\Big)^{2/p}dr.$ (3.21)
Using Hölder inequality and taking $p$-th power,
$\Big|\int_{s}^{t}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}du\Big|^{p}\leq\Big(\int_{s}^{t}|X^{\epsilon}_{s,u}|^{2}du\Big)^{p/2}\Big(\int_{s}^{t}|z^{\epsilon,r}_{u}|^{2}\Big)^{p/2}.$ (3.22)
Applying Lemma 3.2 to $G(u)=|X^{\epsilon}_{s,u}|$ and $q=p/2,$ and then using (3.17),
$\Big\{\mathbb{E}\Big(\int_{s}^{t}|X^{\epsilon}_{s,u}|^{2}du\Big)^{p/2}\Big\}^{2/p}$
$\leq$ $\int_{s}^{t}(\mathbb{E}|X^{\epsilon}_{s,u}|^{p})^{2/p}du$
$\lesssim$ $\int_{s}^{t}|u-s|du$
$=$ $\frac{(t-s)^{2}}{2}.$
Taking ($p/2$)-th power,
$\mathbb{E}\Big(\int_{s}^{t}|X^{\epsilon}_{s,u}|^{2}du\Big)^{p/2}\lesssim|t-s|^{p}.$ (3.23)
On the other hand, from (3.4) we have
$\int_{s}^{t}|z^{\epsilon,r}_{u}|^{2}du\leq\int_{s}^{t}e^{-2\epsilon^{-2}(u-r)}du\leq\frac{\epsilon^{2}}{2}e^{-2\epsilon^{-2}(s-r)}.$
Taking $(p/2)$-th power,
$\Big(\int_{s}^{t}|z^{\epsilon,r}_{u}|^{2}du\Big)^{p/2}\lesssim\epsilon^{p}e^{-p\epsilon^{-2}(s-r)}.$ (3.24)
Combining (3.22), (3.23) and (3.24),
$\mathbb{E}\Big|\int_{s}^{t}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}du\Big|^{p}\leq\epsilon^{p}e^{-p\epsilon^{-2}(s-r)}|t-s|^{p}.$
Taking ($2/p$)-th power and substituting into (3.21),
$\Big\{\mathbb{E}\Big(\int_{0}^{s}\Big|\int_{s}^{t}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}du\Big|^{2}dr\Big)^{p/2}\Big\}^{2/p}$
$\lesssim$ $\int_{0}^{s}\epsilon^{2}e^{-2\epsilon^{-2}(s-r)}|t-s|^{2}dr$
$=$ $\epsilon^{2}(t-s)^{2}\int_{0}^{s}e^{-2\epsilon^{-2}(s-r)}dr$
$\leq$ $\epsilon^{4}(t-s)^{2}.$
Taking ($p/2$)-th power on the above inequality,
$\mathbb{E}\Big(\int_{0}^{s}\Big|\int_{s}^{t}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}du\Big|^{2}dr\Big)^{p/2}\lesssim\epsilon^{2p}(t-s)^{p}.$ (3.25)
We deal with the other term in a similar fashion. Applying Lemma 3.2 to $G(r)=\int_{r}^{t}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}du$ and $q=p/2,$
$\Big\{\mathbb{E}\Big(\int_{s}^{t}\Big|\int_{r}^{t}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}du\Big|^{2}dr\Big)^{p/2}\Big\}^{2/p}\leq\int_{s}^{t}\Big(\mathbb{E}\Big|\int_{r}^{t}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}du\Big|^{p}\Big)^{2/p}dr.$ (3.26)
For fixed $r\in[s,t],$ using Lemma 3.2 to $G(u)=\sqrt{|X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}|},u\in[r,t]$ and $q=p$,
$\Big(\mathbb{E}\Big|\int_{r}^{t}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}du\Big|^{p}\Big)^{1/p}$ (3.27)
$\leq$ $\Big(\mathbb{E}\Big[\int_{r}^{t}|X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}|du\Big]^{p}\Big)^{1/p}$
$\leq$ $\int_{r}^{t}\Big(\mathbb{E}|X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}|^{p})^{1/p}du.$
By (3.4) and (3.17),
$\mathbb{E}|X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}|^{p}\leq e^{-p\epsilon^{-2}(u-r)}\mathbb{E}|X^{\epsilon}_{s,u}|^{p}\leq e^{-p\epsilon^{-2}(u-r)}|u-s|^{p/2}.$
Taking $(1/p)$-th power and substituting it into (3.27),
$\Big(\mathbb{E}\Big|\int_{r}^{t}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}du\Big|^{p}\Big)^{1/p}\leq\int_{r}^{t}e^{-\epsilon^{-2}(u-r)}|u-s|^{1/2}du\leq\epsilon^{2}|t-s|^{1/2}.$ (3.28)
Taking square and substituting it into (3.26),
$\Big\{\mathbb{E}\Big(\int_{s}^{t}\Big|\int_{r}^{t}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}du\Big|^{2}dr\Big)^{p/2}\Big\}^{2/p}\leq\int_{s}^{t}\epsilon^{4}|t-s|dr=\epsilon^{4}|t-s|^{2}.$ (3.29)
Taking $(p/2)$-th power,
$\mathbb{E}\Big(\int_{s}^{t}\Big|\int_{r}^{t}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}du\Big|^{2}dr\Big)^{p/2}\leq\epsilon^{2p}|t-s|^{p}.$ (3.30)
Combining (3.20), (3.25) and (3.30),
$\mathbb{E}|I|^{p}\lesssim|t-s|^{p}.$ (3.31)
It remains to deal with $J$. Writing
$\Big|\int_{0}^{s}\int_{s}^{t}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}F(X^{\epsilon}_{r})dudr\Big|^{p}$ (3.32)
$\lesssim$ $\Big(\int_{0}^{s}\int_{s}^{t}|X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}|dudr\Big)^{p}$
$=$ $\Big[\Big(\int_{0}^{s}\int_{s}^{t}|X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}|dudr\Big)^{2}\Big]^{p/2}$
$\lesssim$ $\Big[\int_{0}^{s}\Big(\int_{s}^{t}|X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}|du\Big)^{2}dr\Big]^{p/2},$
@@
$\Big|\int_{s}^{t}\int_{r}^{t}X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}F(X^{\epsilon}_{r})dudr\Big|^{p}$ (3.33)
$\lesssim$ $\Big(\int_{s}^{t}\int_{r}^{t}|X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}|dudr\Big)^{p}$
$=$ $\Big[\Big(\int_{s}^{t}\int_{r}^{t}|X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}|dudr\Big)^{2}\Big]^{p/2}$
$\lesssim$ $\Big[\int_{s}^{t}\Big(\int_{r}^{t}|X^{\epsilon}_{s,u}z^{\epsilon,r}_{u}|du\Big)^{2}dr\Big]^{p/2},$
so that we notice that the estimate of $J$ is identical to $I$ in view of (3.20), from which the proof is ended. ∎
@@
From Corollary 3.4, we see that $(X^{\epsilon})_{0<\epsilon\leq 1}$ is tight on $C([0,T];\mathbb{R}^{d}),$ which is assumed in [12]. But we do not use tightness throughout this paper.

### v2

@@ 3 Moment Estimates
$=$ $\int_{s}^{t}M^{-1}(X^{\epsilon}_{u})dW_{u}+\int_{s}^{t}M^{-1}(X^{\epsilon}_{u})F(X^{\epsilon}_{u})du-\epsilon\int_{s}^{t}M^{-1}(X^{\epsilon}_{u})dY^{\epsilon}_{u}$
$=:$ $I^{\epsilon,1}(s,t)+I^{\epsilon,2}(s,t)+I^{\epsilon,3}(s,t).$
By Burkholder–Davis–Gundy inequality and ($\mathbf{A4}$),
$\mathbb{E}|I^{\epsilon,1}(s,t)|^{p}\lesssim\mathbb{E}\Big(\int_{s}^{t}|M^{-1}(X^{\epsilon}_{u})|^{2}du\Big)^{p/2}\lesssim(t-s)^{p/2}.$ (3.6)
$\mathbb{E}|I^{\epsilon,2}(s,t)|^{p}\lesssim\mathbb{E}\Big(\int_{s}^{t}|M^{-1}(X^{\epsilon}_{u})F(X^{\epsilon}_{u})|du\Big)^{p}\lesssim(t-s)^{p}\lesssim(t-s)^{p/2}.$ (3.7)
Applying integration-by-part formula and (1.6),
$I^{\epsilon,3}_{i}(s,t)$ (3.8)
$=$ $\epsilon\sum_{j}\int_{s}^{t}M^{-1}_{ij}(X^{\epsilon}_{u})dY^{\epsilon,j}_{u}$
$=$ $\epsilon\sum_{j}[M^{-1}_{ij}(X^{\epsilon}_{t})Y^{\epsilon,j}_{t}-M^{-1}_{ij}(X^{\epsilon}_{s})Y^{\epsilon,j}_{s}]-\epsilon\sum_{j,l}\int_{s}^{t}Y^{\epsilon,j}_{u}dM^{-1}_{ij}(X^{\epsilon}_{u})$
$=$ $\epsilon\sum_{j}[M^{-1}_{ij}(X^{\epsilon}_{t})Y^{\epsilon,j}_{t}-M^{-1}_{ij}(X^{\epsilon}_{s})Y^{\epsilon,j}_{s}]-\epsilon\sum_{j,l}\int_{s}^{t}Y^{\epsilon,j}_{u}\partial_{x_{l}}M^{-1}_{ij}(X^{\epsilon}_{u})dX^{\epsilon,l}_{u}$
$=$ $\epsilon\sum_{j}[M^{-1}_{ij}(X^{\epsilon}_{t})Y^{\epsilon,j}_{t}-M^{-1}_{ij}(X^{\epsilon}_{s})Y^{\epsilon,j}_{s}]-\sum_{j,l}\int_{s}^{t}Y^{\epsilon,j}_{u}Y^{\epsilon,l}_{u}\partial_{x_{l}}M^{-1}_{ij}(X^{\epsilon}_{u})du$
$=:$ $J^{\epsilon,1}_{i}(s,t)+J^{\epsilon,2}_{i}(s,t).$
Since
$J^{\epsilon,1}_{i}(s,t)=\epsilon\sum_{j}[M^{-1}_{ij}(X^{\epsilon}_{t})(Y^{\epsilon,j}_{t}-Y^{\epsilon,j}_{s})+Y^{\epsilon,j}_{s}(M^{-1}_{ij}(X^{\epsilon}_{t})-M^{-1}_{ij}(X^{\epsilon}_{s}))],$
an application of ($\mathbf{A3}$), Proposition 3.1 and the dichotomy assumption yields
$\mathbb{E}|J^{\epsilon,1}_{i}(s,t)|^{p}$ (3.9)
$\lesssim$ $\epsilon^{p}\mathbb{E}(|M^{-1}_{ij}(X^{\epsilon}_{t})|^{p}|Y^{\epsilon,j}_{t}-Y^{\epsilon,j}_{s}|^{p})+\epsilon^{p}\mathbb{E}(|Y^{\epsilon,j}_{s}|^{p}|M^{-1}_{ij}(X^{\epsilon}_{t})-M^{-1}_{ij}(X^{\epsilon}_{s})|^{p})$
$\lesssim$ $\epsilon^{p}\leq(t-s)^{p/2}.$
By Hölder inequality, ($\mathbf{A3}$) and Proposition 3.1,
$\mathbb{E}|J^{\epsilon,2}_{i}(s,t)|^{p}$ $\lesssim$ $\mathbb{E}\Big(\int_{s}^{t}|Y^{\epsilon,j}_{u}Y^{\epsilon,l}_{u}\partial_{x_{l}}M^{-1}_{ij}(X^{\epsilon}_{u})|du\Big)^{p}$ (3.10)
$\lesssim$ $\mathbb{E}\Big(\int_{s}^{t}|Y^{\epsilon,j}_{u}Y^{\epsilon,l}_{u}|du\Big)^{p}$
$\leq$ $(t-s)^{p-1}\int_{s}^{t}\mathbb{E}|Y^{\epsilon,j}_{u}Y^{\epsilon,l}_{u}|^{p}du$
$\lesssim$ $(t-s)^{p}\lesssim(t-s)^{p/2}.$
Combining (3.8)–(3.10),
$\mathbb{E}|I^{\epsilon,3}_{i}(s,t)|^{p}\lesssim(t-s)^{p/2}.$ (3.11)
Combining (3.5)–(3.7) and (3.11), we have
$\mathbb{E}|X^{\epsilon}_{t}-X^{\epsilon}_{s}|^{p}\lesssim(t-s)^{p/2},$ (3.12)
and this finishes the proof of (i). Next we turn to (ii).
Case 1: $\epsilon\geq\sqrt{t-s}$.
From (1.6) we see that
$\mathbb{X}^{\epsilon}_{s,t}=\frac{1}{\epsilon}\int_{s}^{t}X^{\epsilon}_{s,u}\otimes Y^{\epsilon}_{u}du.$
By Hölder inequality, Proposition 3.1 and Proposition 3.2 (i),
$\mathbb{E}|\mathbb{X}^{\epsilon}_{s,t}|^{p}$ $\leq$ $\frac{1}{\epsilon^{p}}(t-s)^{p-1}\int_{s}^{t}\mathbb{E}|X^{\epsilon}_{s,u}\otimes Y^{\epsilon}_{u}|^{p}du$ (3.13)
$\leq$ $\frac{1}{\epsilon^{p}}(t-s)^{p-1}\int_{s}^{t}\sqrt{\mathbb{E}|X^{\epsilon}_{s,u}|^{2p}}\sqrt{\mathbb{E}|Y^{\epsilon}_{u}|^{2p}}du$
$\lesssim$ $\frac{1}{\epsilon^{p}}(t-s)^{p-1}\int_{s}^{t}(u-s)^{p/2}du$
$\lesssim$ $\frac{1}{\epsilon^{p}}(t-s)^{3p/2}\leq(t-s)^{p}.$
Case 2: $\epsilon<\sqrt{t-s}$.
By (3.3),
$\mathbb{X}^{\epsilon}_{s,t}$ (3.14)
$=$ $\int_{s}^{t}X^{\epsilon}_{s,u}\otimes[M^{-1}(X^{\epsilon}_{u})dW_{u}]+\int_{s}^{t}X^{\epsilon}_{s,u}\otimes[M^{-1}(X^{\epsilon}_{u})F(X^{\epsilon}_{u})du]-\epsilon\int_{s}^{t}X^{\epsilon}_{s,u}\otimes[M^{-1}(X^{\epsilon}_{u})dY^{\epsilon}_{u}]$
$=:$ $I^{\epsilon,1}(s,t)+I^{\epsilon,2}(s,t)+I^{\epsilon,3}(s,t).$
By Burkholder–Davis–Gundy inequality, ($\mathbf{A3}$)–($\mathbf{A4}$), Hölder inequality and Proposition 3.2 (i),
$\mathbb{E}|I^{\epsilon,1}_{ij}(s,t)|^{p}$ $=$ $\mathbb{E}\Big|\sum_{k}\int_{s}^{t}X^{\epsilon,i}_{s,u}M^{-1}_{jk}(X^{\epsilon}_{u})dW^{k}_{u}\Big|^{p}$ (3.15)
$\lesssim$ $\sum_{k}\mathbb{E}\Big(\int_{s}^{t}|X^{\epsilon,i}_{s,u}M^{-1}_{jk}(X^{\epsilon}_{u})|^{2}du\Big)^{p/2}$
$\lesssim$ $\mathbb{E}\Big(\int_{s}^{t}|X^{\epsilon,i}_{s,u}|^{2}du\Big)^{p/2}$
$\lesssim$ $(t-s)^{\frac{p}{2}-1}\int_{s}^{t}\mathbb{E}|X^{\epsilon,i}_{s,u}|^{p}du$
$\lesssim$ $(t-s)^{\frac{p}{2}-1}\int_{s}^{t}(u-s)^{p/2}du\lesssim(t-s)^{p},$
@@
$\mathbb{E}|I^{\epsilon,2}_{ij}(s,t)|^{p}$ $=$ $\mathbb{E}\Big|\sum_{k}\int_{s}^{t}X^{\epsilon,i}_{s,u}M^{-1}_{jk}(X^{\epsilon}_{u})F_{k}(X^{\epsilon}_{u})du\Big|^{p}$ (3.16)
$\lesssim$ $\mathbb{E}\Big(\int_{s}^{t}|X^{\epsilon,i}_{s,u}|du\Big)^{p}$
$\leq$ $(t-s)^{p-1}\int_{s}^{t}\mathbb{E}|X^{\epsilon,i}_{s,u}|^{p}du$
$\lesssim$ $(t-s)^{p-1}\int_{s}^{t}(u-s)^{p/2}du$
$\lesssim$ $(t-s)^{3p/2}\lesssim(t-s)^{p}.$
By integration-by-part formula and (1.6),
$I^{\epsilon,3}_{ij}(s,t)$ (3.17)
$=$ $-\epsilon\sum_{k}\int_{s}^{t}X^{\epsilon,i}_{s,u}M^{-1}_{jk}(X^{\epsilon}_{u})dY^{\epsilon,k}_{u}$
$=$ $-\epsilon\sum_{k}\Big\{X^{\epsilon,i}_{s,t}M^{-1}_{jk}(X^{\epsilon}_{t})Y^{\epsilon,k}_{t}-\int_{s}^{t}Y^{\epsilon,k}_{u}d[X^{\epsilon,i}_{s,u}M^{-1}_{jk}(X^{\epsilon}_{u})]\Big\}$
$=$ $-\epsilon\sum_{k}X^{\epsilon,i}_{s,t}M^{-1}_{jk}(X^{\epsilon}_{t})Y^{\epsilon,k}_{t}+\epsilon\sum_{k}\int_{s}^{t}Y^{\epsilon,k}_{u}M^{-1}_{jk}(X^{\epsilon}_{u})dX^{\epsilon,i}_{u}+\epsilon\sum_{k}\int_{s}^{t}Y^{\epsilon,k}_{u}X^{\epsilon,i}_{s,u}d[M^{-1}_{jk}(X^{\epsilon}_{u})]$
$=$ $-\epsilon\sum_{k}X^{\epsilon,i}_{s,t}M^{-1}_{jk}(X^{\epsilon}_{t})Y^{\epsilon,k}_{t}+\epsilon\sum_{k}\int_{s}^{t}Y^{\epsilon,k}_{u}M^{-1}_{jk}(X^{\epsilon}_{u})dX^{\epsilon,i}_{u}+\epsilon\sum_{k,l}\int_{s}^{t}Y^{\epsilon,k}_{u}X^{\epsilon,i}_{s,u}\partial{x_{l}}M^{-1}_{jk}(X^{\epsilon}_{u})dX^{\epsilon,l}_{u}$
$=$ $-\epsilon\sum_{k}X^{\epsilon,i}_{s,t}M^{-1}_{jk}(X^{\epsilon}_{t})Y^{\epsilon,k}_{t}+\sum_{k}\int_{s}^{t}Y^{\epsilon,k}_{u}M^{-1}_{jk}(X^{\epsilon}_{u})Y^{\epsilon,i}_{u}du+\sum_{k,l}\int_{s}^{t}Y^{\epsilon,k}_{u}Y^{\epsilon,l}_{u}X^{\epsilon,i}_{s,u}\partial{x_{l}}M^{-1}_{jk}(X^{\epsilon}_{u})du$
$=:$ $J^{\epsilon,1}_{ij}(s,t)+J^{\epsilon,2}_{ij}(s,t)+J^{\epsilon,3}_{ij}(s,t).$
By Hölder inequality, ($\mathbf{A3}$) and Proposition 3.1,
$\mathbb{E}|J^{\epsilon,1}_{ij}(s,t)|^{p}$ $\lesssim$ $\epsilon^{p}\mathbb{E}(|X^{\epsilon,i}_{s,t}|^{p}|Y^{\epsilon,k}_{t}|^{p})$ (3.18)
$\lesssim$ $\epsilon^{p}\sqrt{\mathbb{E}|X^{\epsilon,i}_{s,t}|^{2p}}\sqrt{\mathbb{E}|Y^{\epsilon,k}_{t}|^{2p}}$
$\lesssim$ $\epsilon^{p}(t-s)^{p/2}\leq(t-s)^{p},$
where we have used the dichotomy assumption in the last inequality.
By ($\mathbf{A3}$), Hölder inequality and Proposition 3.1,
$\mathbb{E}|J^{\epsilon,2}_{ij}(s,t)|^{p}$ $\lesssim$ $\mathbb{E}\Big(\int_{s}^{t}|Y^{\epsilon,k}_{u}Y^{\epsilon,i}_{u}|du\Big)^{p}$ (3.19)
$\leq$ $(t-s)^{p-1}\int_{s}^{t}\mathbb{E}|Y^{\epsilon,k}_{u}Y^{\epsilon,i}_{u}|^{p}du\lesssim(t-s)^{p}.$
By ($\mathbf{A3}$) and Proposition 3.1,
$\mathbb{E}|J^{\epsilon,3}_{ij}(s,t)|^{p}$ $\lesssim$ $\mathbb{E}\Big(\int_{s}^{t}|Y^{\epsilon,k}_{u}Y^{\epsilon,l}_{u}X^{\epsilon,i}_{s,u}|du\Big)^{p}$ (3.20)
$\leq$ $(t-s)^{p-1}\int_{s}^{t}\mathbb{E}|Y^{\epsilon,k}_{u}Y^{\epsilon,l}_{u}X^{\epsilon,i}_{s,u}|^{p}du$
$\leq$ $(t-s)^{p-1}\int_{s}^{t}\sqrt{\mathbb{E}|Y^{\epsilon,k}_{u}Y^{\epsilon,l}_{u}|^{2p}}\sqrt{\mathbb{E}|X^{\epsilon,i}_{s,u}|^{2p}}du$
$\lesssim$ $(t-s)^{p-1}\int_{s}^{t}(u-s)^{p/2}du$
$\lesssim$ $(t-s)^{3p/2}\lesssim(t-s)^{p}.$
Combining (3.17)–(3.20),
$\mathbb{E}|I^{\epsilon,3}_{ij}(s,t)|^{p}\lesssim|t-s|^{p}.$ (3.21)
Combining (3.14)–(3.16) and (3.21),
$\mathbb{E}|\mathbb{X}^{\epsilon}_{s,t}|^{p}\lesssim|t-s|^{p}.$ (3.22)
@@
(i) From Corollary 3.3 and Kolmogorov tightness criterion for rough paths[8, Theorem 3.10], we see that $(\mathbf{X}^{\epsilon})_{0<\epsilon\leq 1}$ is tight in $\mathscr{C}^{0,\alpha}([0,T];\mathbb{R}^{d}).$
[…]
$\sup\limits_{0<\epsilon\leq 1}\mathbb{E}\sup\limits_{0\leq t\leq T}|X^{\epsilon}_{t}|^{p}<\infty,\quad p\geq 1.$ (3.23)
Remark 3.5.
One might expect to show Proposition 3.2 (i) by integrating factor method in classical ODE theory, that is, write
$X^{\epsilon}_{t}=\frac{1}{\epsilon^{2}}\int_{0}^{t}\int_{0}^{s}z^{\epsilon,r}_{s}dW_{r}ds+\frac{1}{\epsilon^{2}}\int_{0}^{t}\int_{0}^{s}z^{\epsilon,r}_{s}F(X^{\epsilon}_{r})drds,$
where $z^{\epsilon,r}_{s}=\phi^{\epsilon}_{s}(\phi^{\epsilon}_{r})^{-1},\kern 5.0pt0\leq r\leq s\leq T$, and $\phi^{\epsilon}$ is the solution of the random ODE
$\dot{\phi}^{\epsilon}_{t}=-\frac{1}{\epsilon^{2}}M(X^{\epsilon}_{t})\phi^{\epsilon}_{t},\kern 5.0pt\phi^{\epsilon}_{0}=\mathrm{I}_{d};$
see [27] for detailed computation. However, this seems incorrect, since the integrand $r\mapsto z^{\epsilon,r}_{s}$ is not $\mathcal{F}_{r}$-adapted, let alone progressively measurable, so the integrand with respect to Brownian motion is not well-defined. This is also pointed out by Blassel [2].